the difference between the ages of two students is 4 years. if the sum of their ages is 425 years. find each students age
Answer by Mathtut(3670)

students a and b. so a-b=4 and a+b=425 solve for a in first equation and you get a=b+4 substitute that into 2nd equation we get (b+4)+b=425 solve for b b= 210.5 so a =214.5......now that is some old students were these bible time students???? lol